OpenStudy (anonymous):

this looks failure simplify 9 ^1/2 (multiplied by) 9^1/2 .....is it 81=/

OpenStudy (anonymous):

answer is 9

OpenStudy (anonymous):

the answer is 9

OpenStudy (anonymous):

arghh im not copying :D

OpenStudy (anonymous):

\[\sqrt{9}\times\sqrt{9}=9\]

OpenStudy (anonymous):

so where does the 1/2 come into play

OpenStudy (anonymous):

exponent of one half means square root

OpenStudy (anonymous):

oh yeah =/ lol

OpenStudy (anonymous):

just like when we wrote in exponential form \[\sqrt{x}=x^{\frac{1}{2}}\]